OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?

zepdrix (zepdrix):

So they gave us the equation of a line in `standard form`: \[\Large\rm 3x-2y=-6\] If we find the slope we can set up the other line. Do you know how to get this line into `slope-intercept` form?

OpenStudy (anonymous):

no not really

zepdrix (zepdrix):

slope-intercept form looks like this \(\Large\rm y=stuff\) Only y on the left side. So to isolate the y, you need to move all that other junk to the right side. Subtracting 3x gives us,\[\Large\rm -2y=-6-3x\]Understand that step?

OpenStudy (anonymous):

mhmm yep

OpenStudy (anonymous):

then your suppose to take 2 away frum both sides to get y by itself?

zepdrix (zepdrix):

Well the negative 2 is `multiplying` y, so we can't take it away by adding/subtracting right?

OpenStudy (anonymous):

nope division

zepdrix (zepdrix):

\[\Large\rm y=\frac{-6-3x}{-2}\]Ok cool. So you'll need to divide `both` the -6 and -3 by -2. Can you simplify that right side?

OpenStudy (anonymous):

yep into \[\frac{ 3x }{ -2 }\] ?

zepdrix (zepdrix):

No, those negatives should cancel out, changing it to positive.

zepdrix (zepdrix):

And don't forget about the other term, the -6 divide -2 gives us 3, yes?

OpenStudy (anonymous):

yeah yeah i 4got about taking out the negatives but thanx u've been a real help ...its kinda hard switching from Geometry back to Algebra 1

zepdrix (zepdrix):

Anyway, this only get's you through half of the problem.\[\Large\rm y=3+\frac{3}{2}x\]So you found that your slope is \(\Large\rm m=\dfrac{3}{2}\) Since the other line is `parallel` to this one, it means they have the `same slope`.

zepdrix (zepdrix):

So our new line will be,\[\Large\rm y=\frac{3}{2}x+b\]It has a different y-intercept than the line we started with, but the same slope (parallel). To find the y-intercept, plug in the coordinate pair they gave you, and solve for b.

OpenStudy (anonymous):

k thnx